OpenStudy (anonymous):

Which of the following is a solution to the equation 1/2(x+4)^1/3=2? a) -3 b) 8 c) 12 d) 60

OpenStudy (anonymous):

60 +4 = 64 64^{1/3}=4 1/2 4 =2 It is d